The area of a rhombus is 7.5 sq. in. Find the length of the two diagonals if one of them is one more than twice the other. Show all work to receive credit.
Answer by scott8148(5884)

area of rhombus is 1/2 the product of the diagonals

let x=diagonal, so 2x+1="other diagonal"

(1/2)(x)(2x+1)=7.5 ... 2x^2+x=15 ... 2x^2+x-15=0 ... (2x-5)(x+3)=0